Back to the NIKA2 main page

Daily Reports

Jan 18th, Tuesday

16:00-21:00; Ioannis: System check begins directly after maintenance around 16:30. NIKA2 acquisition is restarted. After few trials we get a good tuning with 42/5/48 dead kids for Arr1/Arr2/Arr3. The correlated noise in the traces of Arr1 is still visible at 2.6, 5.5 and 8 Hz. We start with very good weather conditions (tau225 ~ 0.15 and baseline variations of 0.5-1 Jy/beam at 2mm) and do pointing and focus on Uranus (beam size 17.7" at 2mm and best focus -1.3mm). Manuel adjusts parameters P2 and P7 of the pointing model and I start a short pointing session with 16 sources that cover an azimuth range of 120-289 degrees and elevation range of 10-73 degrees. The results yield an azimuth correction of 1.81" +/- 1.99" and an elevation correction of -6.29" +/- 1.28". Given the large offset from (0,0) we readjust parameter P2 and run a second short pointing session with 11 sources that cover an azimuth range of 120-283 degrees and elevation range of 27-70 degrees. The final results yield an azimuth correction of 1.13" +/- 0.99" and an elevation correction of 0.47" +/- 1.56" and we decide to keep this settings for the ongoing pool week. The shift is finished with a retuning of NIKA2 in preparation for the night shifts. After few trials we get a good tuning with 38/8/48 dead kids for Arr1/Arr2/Arr3. Throughout the shift tau225 was below 0.2 and the baseline stability was between 0.3 and 4 Jy/beam at 2mm.

21:00-03:00; Matthieu: The night started by calibration on Uranus (pointing, focus, calib1_scan [20220118s193], beammap [20220118s194]) and then a skydip [20220118s195]. Then, WB89-789 (137-21) was observed. The operator (Sergio) mentioned a problem with an antenna temperature control. I redid a focus and it was off by almost 2mm. We managed to get a good focus and started to observe COSMOS (192-16). I did a control pointing and realize that the beam was VERY bad (40 arcsec). The operator then changed the temperature control mode. Around 1h30, I redid a focus (hard to find since it moved again a lot) and started to observe COSMOS. I redid a focus and pointing after 30 min and it did not drift too much. I then started a 1h integration on COSMOS to finish the shift. The weather was excellent (tau~0.15). So, it was very unfortunate to have these technical problems, but Sergio was super efficient to fix it.

Jan 19th, Wednesday

03:00-09:00; Stavroula: I tried to focus the telescope every hour in order to avoid beam distortions. After 1 hour, beam was going towards 20" @2mm and after focusing was dropping at 17.5". NIKA2 tuning done by Ioannis at the end of the shift [IM: good tuning after a few trials with 45/7/57 dead kids for Arr1/Arr2/Arr3]. Besides that, sky conditions were great: tau pretty stable to 0.17-0.18, sky fluctuations between +/- 2Jy/beam for the most of the shift. Concerning the observations: After the end of COSMOS repetition (192-16), I observed the project 160-16: 88min on NGC4321 (currently considered as complete), 87min on NGC3184, 20min on NGC3938. One skydip done roughly at the mid of the shift [20220119s49].

09:00-15:00; Miren & Aishwarya: Opacity was very stable at ~0.18 during the whole shift. However, due to the temperature control problem we had to check pointing and focus very frequently [IM: every 1hr or less]. We started with two repetitions on PSZ2G094 (199-16) and after focus observations were paused for restarting the automatic temperature control [IM: the system is monitored and if any more issues appear we will tackle them on the spot]. Then we observed PSZ2G091 for one repetition (199-16). Switched to 109-21 project and after pointing did 2 reps on L694 (35 and 36/50). Beam was relatively stable, except towards the end of the shift, when it was broader in azimuth (~ 17 x 20 at 2mm). We repeated pointing and focus to improve the beam at the end of the shift.

15:00-21:00; Quang: Observation status: 6/6 hours. Opacity: ~0.2 or less. Maps: 6 maps in NG1333 sources (110-21), 4 maps in B211 (122-16). We did one skydip [20220119s153], 1 calib_1scan [20220119s175], pointings and focuses.

21:00-03:00; Alex: End of scan 179 on B211@122-16. IM retuned the NIKA2 detectors [IM: good tuning after a few trials with 44/8/47 dead kids for Arr1/Arr2/Arr3]. Tau stable around 0.2. Pointing on Uranus, PIIC 2mm trace look strange, might be a cloudsat (alarm went on), first focus offset seems quite large, repeating pointing to validate beamsizes, repeating focus before beammap1scan99sub a [20220119s194], launched a post-focus for verification, beams are stable, waiting 7 min for piic to start the processing of the focus scans... IM restarted the imbfits creation. @calib_1scan [20220119s201] and skydip [20220119s202]. Switch to 143-21, pointing and integration on G202, second scan (205) cancelled due to a temperature control alarm, checked focus still stable; switch to 196-12, integration on COSMOS, tau slowing getting to 0.15, waited 5min45s / 5min55s / 5min25 to get the pointing reduction after the two cosmos maps... no IDL pipeline already opened on mrt-remote:5932 [IM: IDL is always running on desktop #3 of VNC32] -- Waiting time statistic : 01:14 00:48 00:57 00:50 02:44 05:45 5:55 5:25

Jan 20th, Thursday

03:00-09:00; Stavroula: During the observations the beam was pretty stable, except for the first focus check, one hour since I started, where I found it increased and distorted. Sky conditions were great: tau stable to 0.14-0.15 for the most of the shift, gradually rising to 0.18-0.19, sky fluctuations up to +/- 5Jy/beam @ 2mm. Observations: After the end of COSMOS repetition (192-16), I observed the project 160-16: 78min on NGC4536, 115min on NGC3184, 57min on NGC3938. I executed one skydip roughly at 5.30am [20220120s34].

9h00-15h00; Juan: Starting with focus but bad beams until 9h20 UT; Then, manage to focus on Saturn, project 160-16 : 1 rep on NCG7331 , project 199-16: 2 rep PSZ2G094 and 2 rep PSZ2G091. Technical work to check the telescope temperature sensors. Very good weather (tau255 ~ 0.1) at the end of the shift but noise a bit too large. [IM: At the beginning of the shift NIKA2 was tuned and after a few trials we got a good tuning with 47/7/45 dead kids for Arr1/Arr2/Arr3]

15:00-21:00; Luca: We started on time at 15h, the beam size was checked on the secondary calibrators MWC349 and was ~17.3 x 17.5 at 2mm. On the same source, the calib_1scan [20220120s126] has been also performed. After that, we started observing the project 137-21: two sources included in the PI's high priority list have been completed. At 17:45 while doing the focus on MWC349 an elongated beam shape started to appear. More beam measurements and focus on Uranus did not improve the results. Around 18h we performed the skydip [20220120s169]. We resumed focussing and pointing on Uranus but the beam shape continued to be quite elongated, specially at 1.2mm [IM: Beam size was 17.5"x18.7" @ 2mm and 12.5"x17" @ 1mm]. We keep pointing and focussing, waiting for the antenna surface to relax after sunset. This continued until the end of my shift at 21h. [IM: The relaxing period of the antenna was about 1.5 hrs long. During that time and at the beginning of the next shift a focusing cadence of 1hr was implemented].

21:00-03:00; Longji: During the observation, it is difficult to find good beam in array1/3 even at the solution of the focus loop. The beams in array 1/3 have elongated shapes [IM: typically 12"x18"] even after we retuned the KIDs. The shift started with 1 repetition on G202 (143-21) for 1 hour. After that we returned the KIDs [IM: good tuning after a few trials with 34/7/42 dead kids for Arr1/Arr2/Arr3], did 1 calib_1scan [20220120s221] and a beam map [20220120s222] on Uranus and 1 skydip [20220120s223], then we observed COSMOS (192-16) for about 2 hours until the end of the shift.

Jan 21st, Friday

03:00-09:00; Stavroula & Angelos: We couldn't observe with NIKA2. In the beginning of the shift, we checked the focus, and after a few tries, beam remained increased and enlogated, especially at arrays 1 and 3 [IM: ~ 11"x18"]. The problem remained (antenna temperature was normal), so, after some hours of trying [IM: we tried many things, checking the optical path, temperature sensors, retuning NIKA2 etc.], we changed to EMIR at 05:30. We observed the project 170-21: 4 repetitions at 48125-high and 4 repetitions at 48131-high. The problem was affecting EMIR observations too, so we stopped. I wish good luck on solving this!

09:00-15:00; Corentin & Aishwarya: [IM: Observations with NIKA2 restarted at 11:15. NIKA2 was tuned and after a few trials we got a good tuning with 36/6/56 dead kids for Arr1/Arr2/Arr3. Focus on 1633+392. Beam looks much better, 17.3"x17.8" @ 2mm and 11.1"x11.9" @ 1mm. Then calib_1scan on MWC349 (20220121s109), a skydip (20220121s110) and I give the telescope back to the observers.] We did a pointing on 1807+698 which was fine, then 2 reps on PSZ2-091.8. After that we did one pointing on 1807+698 again which was not quite good so we switched to project 109-21. before that we did a pointing on a source near L694-2 but we found it quite weird so we did a pointing and 2 focus on 1633+382 (the second focus seemed quite good). Finally we did one last pointing on this source.

15:00-21:00; Luca: The shift started at 15h, the focus was recent (14:40) and the beam size was good. We moved soon to the 137-21 targets using MWC349 as pointing source. Two objects were observed until 17:45 with no issues except some noise instabilities (deviations up to +/- 10 mJy/beam at 2 mm). At 18:15 the skydip [20220121s177] was performed followed by the calib_1scan [20220121s184] on Uranus that was at ~ 70ยบ elevation. At ca. 19h I resumed the science observation (137-21) after doing a focus on the nearby QSO 0355+508. The beam was OK both at 2mm and 1.2mm, very symmetric. In this second half of the shift, the atmosphere was much more stable with baseline deviations < 1mJy/beam. Three sources of the high priority list of 137-21 have been observed today during the shift.

21:00-03:00; Longji: [IM: At the beginning of the shift NIKA2 was tuned and after a few trials we got a good tuning with 41/9/46 dead kids for Arr1/Arr2/Arr3. Also Manuel changed parameter P7 of the (NIKA2) pointing model, in an attempt to move the pointing offsets cloud closer to (0",0"). For the last two days it was centered around (0",5"). Will check again the pointing offsets cloud in a couple of days to see the effect of the P7 adjustment. IDL license is occupied and sent e-mail to all NIKA2 observers in case someone is using it.] We retuned the KIDs and checked the focus at the beginning of the shift and everything was fine. Then we did 1 calib_1scan [20220121s213] and a beam map [20220121s215] on Uranus, and observed WB89-501 (137-21) for 1.2h from 22:30. After that, we made 1 skydip [20220121s254] and observed COSMOS (192-16) in the rest of the time. The focus and beam were stable during the shift.

Jan 22nd, Saturday

03:00-09:00; Stavroula & Angelos: The focus was stable, beam didn't increased significantly during the observations. Sky conditions were great: tau stable to 0.12-0.14, sky fluctuations up to +/-2Jy/beam @ 2mm. Observations: After the end of COSMOS repetition (project 192-16), we observed the project 160-16: 78min on NGC4536, 87min on NGC3184 (currently considered as complete), 20min on NGC3938. Also, we observed the project 199-16: 50min on PSZ2-G160.8. We executed one skydip [20220122s50] during the shift.

09:00-15:00; Miren: [IM: At the beginning of the shift NIKA2 was tuned and after a few trials we got a good tuning with 37/5/46 dead kids for Arr1/Arr2/Arr3. It has been seen systematically this week (and in the past?) that the tuning works very well (delta function centered at 0) almost always after we move the telescope to a new position with "track; start", but I am not sure why.] Took over at 8h25 UT and started with a pointing, focus and calib_1scan [20220122s67] on MWC349. Project 199-16: 2 reps on PSZ2G094 (18/24), 1 rep on PSZ2G084 (16/22) and 2 reps on PSZ2G081 (24/44). A focus at 10h40 UT. Sky fluctuations up to +/- 5Jy/beam @ 2mm. Then moved to project 109-21 and did 2 repetitions on L694 (37 and 38/50), pointing on 1749+096. Focus at 12h30 UT. Last pointings showed a very elongated beam at 1mm (20x11) combined with increasing sky fluctuations, so decided to stop observing 109-21. Several tries to correct from pointing and focus at the end of the shift. Opacity was stable during the shift from 0.13 to 0.15. Also made a skydip [20220122s68].

15:00-21:00; Quang & Luca: (A) Quang, 3pm-5.45pm. Observation status: 2.45/2.45 hours. Opacity: 0.1-0.25. Maps: 4 maps in B2111(122-16). We did one skydip [20220122s125], pointings and focuses. (B) From 17h to 21h Luca took over the observations and he moved to the project 137-21. Around 17:30 calib_1scan was done on Uranus [20220122s152]. Two targets included in the high priority list the 137-21 project were observed. The beam size was good throughout the shift (see notes in nikaw-21 TAPAS).

21:00 - 03:00: Nico and Ioannis waited for the previous script completion, then tuning, focus and pointing on Uranus, then beammap on Uranus (~22h15 LT) [20220122s205], calib_1scan [20220122s206]. 1 rep of 143-21/observer_G202. Focus, pointing, then 192-16/observe_cosmos from 12pm (LT). focus and pointing on 0823+033 around 2h20 LT. Back to Cosmos for the remaining time of this shift.

Jan 23rd, Sunday

03:00-09:00; Stavroula: The focus was stable, beam didn't increased significantly during the observations. Sky conditions: tau 0.2-0.3, steady atmosphere and sky fluctuations +/-2Jy/beam up to +/-5Jy/beam @ 2mm. Observations: After the end of COSMOS repetition (project 192-16), I observed the project 160-16: 116min on NGC4536. After 6.30am, the opacity was rising and I kept going with auxiliary scans of mapping the atmosphere conditions, as the situation seemed temporary. Two skydip scans were executed in tau~0.3-0.35 [20220123s40 and 20220123s54]. Good luck to the rest of the day!

09:00-15:00; Antonio & Marco: After the tuning of NIKA2 by IM [IM: After a few trials we got a good tuning with 64/7/67 dead kids for Arr1/Arr2/Arr3.], we took over at 9h48 with tau225 around 0.3 and started with Project nikaw-21: pointing and focus on MWC349 -> large beams (18.6" and 14"), we move towards NGC7027 (19" and >14"). Tau around 0.33, sky fluctuations up to +/- 10Jy/beam @ 2mm. Move to Project 137-21 at 10:50. Pointing, focus, pointing on 2013+370 and then on the last high priority source 19383+2711. Only half of the rounds done because the source is too high. Move to the low priority source WB89-083. Done. Pointing, focus but sky fluctuations up to +/- 20Jy/beam @ 2mm, impossible to estimate the focus ... the shift ends with skydip [20220123s122] and then we move on to EMIR [IM: According to the weather prediction the opacity and atmosphere fluctuations are expected to increase over the next hours. Changing to EMIR backup observations and project 174-21]. At the end of the shift the opacity oscillates up to 0.45.

15:00-21:00; Ioannis: No NIKA2 observations due to bad weather. At 15:00 we changed to EMIR backup projects but the weather was too bad, even for 3mm observations, so we stopped. Weather degraded further during the shift.

21:00-03:00; Ioannis: Bad weather and some snow until 01:30. Tau225 oscillations between 0.4 and 0.6. Around 01:30 conditions improve and we start with EMIR backup observations and project 170-21.

Jan 24th, Monday

03:00-09:00; Golshan: We continued observing with EMIR since the weather was not good enough for NIKA2. We observed project 170-21 and launched 4 repetitions on '32522-high' for an hour. Then corrected the pointing and focus and continued with 4 repetitions of '32597-high' for another hour. After correcting the pointing again, we launched two more scripts on the same source. Afterward, the observations were stopped due to snow and unstable weather. We checked the pointing after about an hour, but the weather was not good enough to observe to the end of the shift.

09:00-15:00; Ioannis: Bad weather, high opacity and atmospheric instability (tau225 ~ 0.4-0.8) throughout the shift. Continued with EMIR setup and checking with pointing on bright sources every ~ 30 mins to see if conditions improve.

15:00-21:00; Ioannis: Around 17:00 opacity began to stabilize around 0.4-0.5. Restart EMIR backup observations (project 174-21) until the end of the shift.

21:00-03:00; Ioannis: Continued with EMIR observations due to high opacity (~0.4). Observed sources 02524 (5 reprtitions) and 09539 (1 repetition) of project 174-21. Slow opacity improvement towards the end of the shift.

Jan 25th, Tuesday

03:00-10:30; Stavroula & Ioannis: Opacity oscillated between 0.35 and 0.5, so we continued with EMIR back up observations until the start of the telescope maintenance. Observed sources 09530 (3 reprtitions) and 14059 (2 repetitions) of project 174-21.

DailyReportsNika2Pool180122 (last edited 2022-01-25 08:16:27 by NikaBolometer)